What is the solution of the system?
x=-2y
x+6y=8

Answer by EdenWolf(517)

We can solve this system by using the substitution method.
x=-2y
x+6y=8
We know that x = -2y, so we can plug in -2y into the second equation instead of X.
-2y+6y=8
Now solve for y.
-2y+6y=8
4y=8
y=2
Now that we know what y is, we can plug in 2 into X in any of the two equations. We can plug Y into the first equation as 2.
x = -2(2)
x = -4
ANSWER: x = -4, y = 2. (-4,2)